# Improved Contact Graph Routing in Delay Tolerant Networks with Capacity and Buffer Constraints

### v2

@@ IV-C Booking resources for forwarding safety margin
IV-D Information sharing in the network
@@ V Benchmark: source-routing CGR
Then, the source node performs a route search using Yen’s algorithm with Dijkstra searches. A list of the $k$ (Yen’s constant) shortest routes is generated. This list is kept until a change occurs in the CP, such as the end of a contact. In this case, the route list is discarded and a new list is generated, as in route-list pruning [28]. A list of candidate routes is constructed as explained in Section II-C3. When no candidate route is found in this list, the bundle is delayed and the route search is performed again after one time step.
@@ VI-A Environment
The CP is constructed for a 24-hour scenario starting from 8:55 pm on August 19, 2020. We assume that all the contacts have the same rate $R_{\text{C}_{j}}$.
A snapshot of the GSs and one Walker constellation with $N_{\text{sat}}=16$ at 11:18:08 pm on August 19, 2020 is illustrated in Figure 7. It was generated using MATLAB Satellite Communications Toolbox [38]. Using cislunar constellations to evaluate the algorithm will be considered in future work (see also the conclusions).
@@ VI-C Other simulation parameters and assumptions
Yen’s constant $k$ is chosen equal to $10$. Finally, we assume that there is no external disruption in the network. That is, the provided CP is valid and no rerouting is needed on intermediate nodes due to external factors. Given these assumptions, there is therefore no need to consider any safety margin for contact capacity and buffer space (Section IV-C).
@@ VII-A Impact of contact-capacity management
In this section, we only evaluate the performance of the proposed contact-capacity management. Therefore, we do not consider the constraints associated with buffer limits, i.e., the maximum node buffer capacity $B_{\text{max}}=~{}\infty$ for all nodes.
Figure 8 shows the average time spent in the network for $N_{\text{sat}}=16$ (constellation $3$) as function of the number of bundles generated at source $N_{\text{b}}$. Three values of contact data rate $R_{\text{C}_{j}}$ are considered: $200$, $400$, and $800$ bps for all contacts. As expected, one can see that the time spent in the network increases with the number of bundles and with decreasing data rate. With low load (e.g., $N_{\text{b}}<400$ for $400$ bps), the proposed algorithm and the SR-CGR have similar performance. As $N_{\text{b}}$ increases, the gap between the curves increases, with less time spent in the network when using our proposed method. This is due to the increased number of rerouting events during forwarding time when using SR-CGR. This number of rerouting events is shown in Figure 9 for the same $N_{\text{sat}}$ (constellation $3$) and $R_{\text{C}_{j}}$ values as function of $N_{\text{b}}$. With low load, there are only few overbooking events happening on the contacts. In this case, traditional SR-CGR has satisfactory performance. When the number of bundles increases, the overbooking phenomenon increases, making rerouting more likely. In case of such overbooking events, intermediate nodes try to overcome this issue by rerouting the bundle from their location toward the destination. Such new routes are longer than the ones avoiding overbooking, computed using the proposed algorithm. Moreover, if no route is found from the node, the bundle has to be rerouted from the source (see the assumptions on the benchmark) also causing additional delay. The proposed algorithm is more efficient because it computes the shortest route directly from the source node while avoiding overbooking. Indeed, as expected, our proposed solution required no rerouting due to contact capacity check failures during the simulation.
@@
Then, Figure 10 illustrates the time spent in the network with $4$, $8$, and $16$ satellites (Constellations $1$, $2$, and $3$, respectively). The contact data rate is $400$ bps. With four satellites (one satellite per orbital plane), the number of nodes is limited and they are more distant compared to constellations with higher numbers of satellites. Thus, the number of communication opportunities is limited, which constrains the routing degrees of freedom. In this case, the proposed algorithm shows a small improvement compared to the benchmark as seen in Figure 10(a). When the number of satellites increases to $8$ or $16$, our algorithm shows greater improvements compared to the benchmark, particularly with high load ($N_{\text{b}}>1000$ in figures 10(b) and 10(c)).
@@
Figure 10: Average time spent in the network ($R_{\text{C}_{j}}=400$ bps).
Figure 11 illustrates the evolution of the CP size over time, to assess the complexity induced by the proposed algorithm. The Y-axis shows the number of bundles in the CP normalized by its initial size, which is of $3478$ contacts for Constellation $3$ (see Table III). The baseline curve, without capacity management, depicts the evolution of the CP size over time as contacts expire and are removed from the CP. Newly appearing contacts are not counted in this scenario. The other curves represent the evolution of the CP size with our capacity management solution. In this case, splitting contacts can either increase the CP size or leave it unchanged if contacts are shortened. Expired contacts are still removed from the CP. We observe that our capacity management has a negligible impact on the CP size and route search complexity under low load conditions. For instance, with $N_{\text{b}}=200$, the maximum difference between the baseline and our split CP size is $1\%$ and $0.3\%$ for $R_{\text{C}_{j}}=200$ and $800$ bps, respectively. Under high load conditions, the difference becomes more significant. For example, with $N_{\text{b}}=2000$, the increased frequency of bundle generation and routing forces the use of late contacts, as early contacts become fully booked. Consequently, the CP size remains elevated until both split and original contacts expire. The highest difference with $N_{\text{b}}=2000$, observed for $R_{\text{C}_{j}}=200$ at $t=1920$, is approximately $10\%$ compared to the baseline.
@@ VII-B Impact of buffer management
In this section, we evaluate the performance of the proposed buffer management solution. No contact capacity restriction is implemented, i.e., $R_{\text{C}_{j}}=\infty,\;\forall j$.
@@
Figure 14 illustrates the performance impact of constellations with different sizes ($1$, $2$, and $3$) under buffer constraints, using $N_{\text{b}}=400$. In networks with fewer nodes, each node experiences higher utilization. With $N_{\text{sat}}=4$, the limited number of communication opportunities results in similar performance between both solutions, as the additional delay from limited storage capacity affects both approaches equally. Increasing the number of satellites to $8$ highlights the effectiveness of our solution, which results in less time spent in the network than the benchmark. Additionally, when the constellation is expanded to $16$ satellites, our proposed solution achieves very low delay, even with low buffer capacity per node.
@@
VIII Conclusions
@@
To manage contact capacity, we introduced contact splitting. It enables tracking of contact portions previously consumed by routed bundles. For the buffer management, we proposed two techniques: temporary bundle-specific contact splitting and edge pruning. These techniques help avoid any portion or succession of contacts that might result in buffer overflows. These temporary modifications depend on the forecast buffer tables that track the future states of the node buffer occupancy levels as a function of time.
@@
Limitations. Although the proposed solution offers improvements over the state-of-the-art while meeting capacity and buffer constraints, our solution remains centralized. The examination of an efficient distributed solution is therefore an interesting research direction for improvement of this work.
@@
Future work. We plan to assess proposed cislunar constellations, from a communication perspective, using the presented routing technology. We believe that the proposed routing algorithm may help with the design of the communication and storage systems to be embedded in the satellites of such lunar constellations.
@@
We also plan to investigate traffic-level solutions rather than a per-bundle solution. In the state-of-the-art, like in this work, the route is chosen based on optimal criteria at the bundle level, not for the entire traffic to be transmitted. Therefore, this per-bundle strategy is unlikely to be optimal for the overall communication performance. Finally, as mentioned in the previous “limitations” subsection, investigating distributed solutions rather than a centralized one is also of interest.

### v3

@@ IV-D Booking resources for forwarding safety margin
IV-E Information sharing in the network
@@ V Benchmark: source-routing CGR
To our knowledge, no existing benchmark was available under the centralized, perfect-CP assumptions we consider. We therefore describe our source-routing CGR benchmark in this section.
@@
Then, the source node performs a route search using Yen’s algorithm with Dijkstra searches. A list of the $k$ (Yen’s constant) shortest routes is generated. This list is kept until a change occurs in the CP, such as the end of a contact. In this case, the route list is discarded and a new list is generated, as in route-list pruning [49]. A list of candidate routes is constructed as explained in Section II-D3. When no candidate route is found in this list, the bundle is delayed and the route search is performed again after one time step.
@@ VI-A Environment
The CP is constructed for a 24-hour scenario starting from 8:55 pm on August 19, 2020. We assume that all the contacts have the same rate $R_{j}$.
A snapshot of the GSs and one Walker constellation with $N_{\text{sat}}=16$ at 11:18:08 pm on August 19, 2020 is illustrated in Figure 7. It was generated using MATLAB Satellite Communications Toolbox [55]. Using cislunar constellations to evaluate the algorithm will be considered in future work (see also the conclusion).
@@ VI-C Other simulation parameters and assumptions
Yen’s constant $k$ is chosen equal to $10$. Finally, we assume that there is no external disruption in the network. That is, the provided CP is valid and no rerouting is needed on intermediate nodes due to external factors. Given these assumptions, there is therefore no need to consider any safety margin for contact capacity and buffer space (Section IV-D).
@@ VII Simulation results
In our idealistic simulation setting, the delivery ratio is one: all bundles eventually reach the destination as the simulation runs until all bundles are delivered (possibly after reroutes, as described in Section V for the benchmark). The impact of buffer size on performance (and thus on buffer utilization) is illustrated in the Figure 12 where buffer size is varied.
@@ VII-A Impact of contact-capacity management
In this section, we only evaluate the performance of the proposed contact-capacity management. Therefore, we do not consider the constraints associated with buffer limits, i.e., the maximum node buffer capacity $B_{\text{max}}=$$\infty$ for all nodes.
Figure 8 shows the average time spent in the network for $N_{\text{sat}}=16$ (constellation $3$) as function of the number of bundles generated at source $N_{\text{b}}$. Three values of contact data rate $R_{j}$ are considered: $200$, $400$, and $800$ bps for all contacts. As expected, one can see that the time spent in the network increases with the number of bundles and with decreasing data rate. With low load (e.g., $N_{\text{b}}<400$ for $400$ bps), the proposed algorithm and the SR-CGR have similar performance. As $N_{\text{b}}$ increases, the gap between the curves increases, with less time spent in the network when using our proposed method. This is due to the increased number of rerouting events during forwarding time when using SR-CGR. This number of rerouting events is shown in Figure 9 for the same $N_{\text{sat}}$ (constellation $3$) and $R_{j}$ values as function of $N_{\text{b}}$. With low load, there are only few overbooking events happening on the contacts. In this case, traditional SR-CGR has satisfactory performance. When the number of bundles increases, the overbooking phenomenon increases, making rerouting more likely. In case of such overbooking events, intermediate nodes try to overcome this issue by rerouting the bundle from their location toward the destination. Such new routes are longer than the ones avoiding overbooking, computed using the proposed algorithm. Moreover, if no route is found from the node, the bundle has to be rerouted from the source (see the assumptions on the benchmark) also causing additional delay. The proposed algorithm is more efficient because it computes the shortest route directly from the source node while avoiding overbooking. Indeed, as expected, our proposed solution required no rerouting due to contact capacity check failures during the simulation.
@@
Then, Figure 10 illustrates the time spent in the network with $4$, 88, and $16$ satellites (Constellations $1$, 22, and $3$, respectively). The contact data rate is $400$ bps. With four satellites (one satellite per orbital plane), the number of nodes is limited and they are more distant compared to constellations with higher numbers of satellites. Thus, the number of communication opportunities is limited, which constrains the routing degrees of freedom. In this case, the proposed algorithm shows a small improvement compared to the benchmark as seen in Figure 10(a). When the number of satellites increases to $8$ or $16$, our algorithm shows greater improvements compared to the benchmark, particularly with high load ($N_{\text{b}}>1000$ in figures 10(b) and 10(c)).
@@
Figure 10: Average time spent in the network ($R_{j}=400$ bps).
Figure 11 illustrates the evolution of the CP size over time, to assess the complexity induced by the proposed algorithm. The Y-axis shows the number of bundles in the CP normalized by its initial size, which is of $3478$ contacts for Constellation $3$ (see Table III). The baseline curve, without capacity management, depicts the evolution of the CP size over time as contacts expire and are removed from the CP. Newly appearing contacts are not counted in this scenario. The other curves represent the evolution of the CP size with our capacity management solution. In this case, splitting contacts can either increase the CP size or leave it unchanged if contacts are shortened. Expired contacts are still removed from the CP. We observe that our capacity management has a negligible impact on the CP size and route search complexity under low load conditions. For instance, with $N_{\text{b}}=200$, the maximum difference between the baseline and our split CP size is $1\%$ and $0.3\%$ for $R_{j}=200$ and $800$ bps, respectively. Under high load conditions, the difference becomes more significant. For example, with $N_{\text{b}}=2000$, the increased frequency of bundle generation and routing forces the use of late contacts, as early contacts become fully booked. Consequently, the CP size remains elevated until both split and original contacts expire. The highest difference with $N_{\text{b}}=2000$, observed for $R_{j}=200$ at $t=1920$, is approximately $10\%$ compared to the baseline.
@@ VII-B Impact of buffer management
In this section, we evaluate the performance of the proposed buffer management solution. No contact capacity restriction is implemented, i.e., $R_{j}=\infty,\;\forall j$.
@@
Figure 14 illustrates the performance impact of constellations with different sizes ($1$, 22, and $3$) under buffer constraints, using $N_{\text{b}}=400$. In networks with fewer nodes, each node experiences higher utilization. With $N_{\text{sat}}=4$, the limited number of communication opportunities results in similar performance between both solutions, as the additional delay from limited storage capacity affects both approaches equally. Increasing the number of satellites to $8$ highlights the effectiveness of our solution, which results in less time spent in the network than the benchmark. Additionally, when the constellation is expanded to $16$ satellites, our proposed solution achieves very low delay, even with low buffer capacity per node.
@@
VIII Future work: traffic-level extension and information sharing
A natural extension is to move from per-bundle routing to traffic-level optimization with priorities. We maintain two priority-specific contact plans (CPs): High Priority (HP), non-preemptible, and Low Priority (LP), preemptible. Scheduling HP traffic consults the HP CP and writes the reservation into both CPs. Any overlapping LP reservations are de-scheduled, their volumes returned to the LP CP, adjacent segments with $t_{\mathrm{end}}=t_{\mathrm{start}}$ are merged, and affected LP bundles are re-routed. Scheduling LP traffic consults and updates only the LP CP. Its reservations may later be invalidated by HP bookings. Regarding information sharing in the network: the idea is then to keep the HP CP globally consistent, allow the LP CP to be useful yet preemptible, and bound control overhead while preserving the feasibility guarantees of the underlying CGR/SABR search. Other future directions (e.g., cislunar evaluation, distributed solutions, and the limitations listed above) are outlined in the conclusion (Section IX).
IX Conclusion
@@
To manage contact capacity, we introduced contact splitting: it turns the initial contact set $V_{c}^{0}$ into the set $V_{c}$ of split contacts and enables tracking of contact portions previously consumed by routed bundles. For the buffer management, we proposed two techniques: temporary bundle-specific contact splitting (again modifying $V_{c}$) and edge pruning (removing from $E_{c}^{0}$ the edges that would cause overflow). These techniques help avoid any portion or succession of contacts that might result in buffer overflows as the route search runs on the feasible contact graph $(V_{c},E_{c})$. These temporary modifications depend on the forecast buffer tables that track the future states of the node buffer occupancy levels as a function of time.
@@
Limitations. Although the proposed solution offers improvements over the state-of-the-art while meeting capacity and buffer constraints, our solution remains centralized. Our evaluation assumes perfect contact-plan knowledge and no random disruptions. The impact of imperfect or delayed information is not assessed. We do not evaluate the algorithm under high error rates, random contact disruptions, or bursty traffic. The effectiveness under stochastic contact or buffer variations is left for future work. Quantitative evaluation of computational overhead for large-scale networks (e.g. many more satellites or longer time horizons) is also left for future work. We do not evaluate the overhead of global information exchange (e.g. forecast buffer tables, CP updates) or its feasibility in bandwidth-constrained deployments. Architectures such as Time-Sensitive Networking (TSN) may offer useful references for such analysis in future work. Validation on platforms such as the ONE Simulator or ION could be considered in future work to enhance reproducibility. The examination of an efficient distributed solution is therefore an interesting research direction for improvement of this work.
@@
Future work. One could assess proposed cislunar constellations, from a communication perspective, using the presented routing technology. We believe that the proposed routing algorithm may help with the design of the communication and storage systems to be embedded in the satellites of such lunar constellations.
@@
One could also investigate traffic-level solutions rather than a per-bundle solution (see Section VIII for a concrete extension with priority-based CPs and information sharing). In the state-of-the-art, like in this work, the route is chosen based on optimal criteria at the bundle level, not for the entire traffic to be transmitted. Therefore, this per-bundle strategy is unlikely to be optimal for the overall communication performance. Evaluating network throughput (e.g., delivery throughput) could be considered in the scope of this future work. Finally, as mentioned in the previous “limitations” subsection, investigating distributed solutions rather than a centralized one is also of interest. As noted in the limitations, future work may also include the directions listed there (ONE/ION validation, disruptions or imperfect CP, information-exchange overhead).
Appendix A Summary of Symbols
Table IV summarizes the mathematical notation.
TABLE IV: Mathematical notation.
Symbol Meaning
$\mathcal{N}$ Set of nodes
$V_{c}^{0}$ (or CP), $V_{c}$ $V_{c}^{0}$ = set of contacts in the contact plan (before splitting), $V_{c}$ = set of split contacts ($V$ is for vertices in the contact graph)
$\mathcal{P}_{f}(S,t_{0})$ Set of feasible routes for bundle of size $S$ and release time $t_{0}$ (paths in $(V_{c}^{0},E_{c}^{0})$ that are schedulable, contact-capacity feasible, and buffer feasible)
$E_{c}^{0}$, $E_{c}$ Set of edges: $E_{c}^{0}$ = connectivity-only ($(i\to j)$ with $v_{i}=u_{j}$), $E_{c}$ = set of feasible edges (after pruning; $E_{c}\subseteq E_{c}^{0}$)
$\text{C}_{i}$ Contact $\text{C}_{i}=(u_{i},v_{i},t_{1,i},t_{2,i},R_{i},\delta_{i})$
$\text{C}_{s}$, $\text{C}_{\text{end}}$ (CRP) Selected contact at current iteration; final contact of best route found (receiver = destination)
$t_{1,i}$, $t_{2,i}$ Start and end time of contact $\text{C}_{i}$
$t_{e_{1}}$, $t_{e_{2}}$ Contact splitting: boundaries of the removed segment ($t_{1,j}\leq t_{e_{1}}\leq t_{e_{2}}\leq t_{2,j}$; interval $(t_{e_{1}},t_{e_{2}})$ is erased)
$R_{i}$, $\delta_{i}$ Data rate and one-way propagation delay of contact $\text{C}_{i}$ (bps)
$u_{i}$, $v_{i}$ Sender and receiver nodes of contact $\text{C}_{i}$
$s$, $d$ Source and destination nodes of a bundle
$t_{0}$ Bundle release time at source
$S_{i}$ Bundle size (bits); $S_{i}$ = size of bundle $i$
$b$ Bundle being routed (buffer management, Algorithm 2)
$P$, $P_{f}$ Route (path) and feasible route; $P=(\text{C}_{1},\ldots,\text{C}_{k})$. Subscript $i_{m}$ = contact index along path (vertex $m$)
$P^{*}$ Optimal route (minimum arrival time)
$t_{\text{Tx,}i}^{\text{C}_{j}}$ Transmission time of bundle $i$ on contact $\text{C}_{j}$
$\tau_{i}$ Transmission start time on contact $\text{C}_{i}$
$f_{i}$ Arrival (finish) time at the receiver $v_{i}$ of contact $\text{C}_{i}$; $f_{i}=\tau_{i}+S/R_{i}+\delta_{i}$. Subscript $s$ for selected contact $\text{C}_{s}$ (CRP)
$A(P)$ (Estimated) Arrival time at destination for route $P$ with a given contact graph
$W_{i}$ Feasible start-time window for contact $\text{C}_{i}$: $W_{i}=[t_{1,i},t_{2,i}-S/R_{i}]$
Symbol Meaning
$n$ Node (generic index; e.g. in $B_{n}(t)$, $\bar{B}_{n}(t)$)
$I_{n}$ Time intervals the bundle is stored in the buffer at node $n$
$B_{n}(t)$, $B_{\text{max}}$ Buffer capacity at node $n$, time $t$; in this paper $B_{n}(t)=B_{\text{max}}$ (constant). $B_{\text{max}}$ = maximum buffer capacity (bits).
$Q_{n}(t)$ Booked buffer load at node $n$, time $t$ (due to previously scheduled bundles)
$\bar{B}_{n}(t)$ Residual buffer capacity at node $n$, time $t$: $\bar{B}_{n}(t)=\max\{0,B_{n}(t)-Q_{n}(t)\}$
$\mathcal{Z}_{\text{b}}$ Set of nodes with potential buffer overflow for bundle $b$
$Z_{k}$ Node with potential overflow ($Z_{k}\in\mathcal{Z}_{\text{b}}$)
$\mathcal{O}_{Z_{k}}$ Set of potential overflows at node $Z_{k}$
$t_{1,o_{l}}$, $t_{2,o_{l}}$ Start and end time of overflow $o_{l}$
$T_{1,s}$, $T_{2,s}$ (CRP) Sets of overflow start and end times at the receiver of $\text{C}_{s}$ (used in buffer-aware route search)
$(V_{c}^{0},E_{c}^{0})$, $(V_{c},E_{c}^{0})$ Connectivity-only contact graph (before and after splitting)
$(V_{c},E_{c})$ Feasible contact graph (route search runs on this graph)
$N_{\text{sat}}$ Number of satellites in the constellation
$N_{\text{p}}$ Number of orbital planes
$N_{\text{b}}$ Number of bundles (traffic load)
$\Delta\Omega$ Angular separation between adjacent orbital planes
$k$ Number of contacts in a path $P=(\text{C}_{1},\ldots,\text{C}_{k})$; or number of shortest routes in Yen’s algorithm (implementation)
FEAP-CB Feasible Earliest-Arrival Path with Capacity and Buffer constraints